# Primordial magnetogenesis from a supercooled dynamical electroweak phase transition

### v1

@@ 2.1 Magnetic fields at production
where $\Delta V_{H}=m_{h}^{4}/16\lambda_{H}$ with $m_{h}\approx 125\,{\rm GeV}$ and $\lambda_{H}\approx 0.126$ is the energy released by the Higgs potential only, $V_{H}=-m_{h}^{2}h^{2}/4+\lambda_{H}h^{4}/4$; [^1] In theories with an extra singlet scalar $s$, the total energy released can be decomposed as $\Delta V=\Delta V_{s}+\Delta V_{H}$. The energy $\Delta V_{s}$ released from the singlet sector does not source visible magnetic fields. In some models, the singlet is charged under a hidden gauge symmetry, in which case $\Delta V_{s}$ may generate hidden magnetic fields. However, even in the presence of vector kinetic mixing, the contribution of hidden magnetic fields to the visible sector remains negligible [35]. $\epsilon\sim 0.1$ is an efficiency factor for producing magnetic fields [36]; $\kappa$ is given by a sum over physical process leading to magnetic fields. For super-cooled phase transitions with $\alpha\gg 1$, the dominant contribution comes from bubble collisions [^2] Another contribution arises from the detonating expansion of an ultra-relativistic bubble (with velocity $v_{\rm wall}\simeq 1$ for supercooled phase transitions), that transfers to the plasma a fraction $\kappa_{\rm exp}=\frac{\alpha(1-\kappa_{\rm col})^{2}}{0.73+0.083\sqrt{\alpha(1-\kappa_{\rm col})}+\alpha(1-\kappa_{\rm col})}$ (2.2) of the potential energy [37, 18]. The $1-\kappa_{\rm col}$ suppression makes this contribution to $\kappa=\kappa_{\rm col}+\kappa_{\rm exp}$ negligible, at least unless $\alpha$ is huge [18].
@@ 2.2 Magnetic fields today
Instead, during the initial radiation-dominated era, magnetic fields are coupled to the cosmological plasma and undergo inverse cascades. Neglecting the expansion of the universe, the coherence length grows as $\lambda\propto\tau^{p_{\lambda}}$, while the magnetic field strength decreases as $B\propto\tau^{-p_{B}/2}$, where $\tau$ denotes the proper time. Various estimates for the two exponents exist in the literature [41, 39, 42, 40], mainly depending on whether a fraction of the magnetic fields has helical shape, as magnetic helicity $\propto B^{2}\lambda$ is conserved in a plasma with high conductivity:  [^4] Magnetic helicity density $\mathbf{A}\cdot\mathbf{B}$ where $\mathbf{B}=\nabla\times\mathbf{A}$ is a measure of the twisting, linking, and knottedness of magnetic field lines. A helical field increases its correlation length as magnetic energy dissipates, driving an inverse cascade of energy from small to large scales.
@@
This means that even a subdominant helical component of the magnetic fields tends to become dominant at late times. Inverse cascades can also arise in the non-helical case, since bubble collisions and parity-violating processes during the phase transition can induce helicity in the plasma velocity field. Taking into account the expansion of the universe — where the scale factor evolves as $a\propto\tau$ during the radiation-dominated era — yields
@@
$B_{0}(\lambda_{0})\approx\begin{cases}8~{}10^{-12}\,{\rm Gauss}\,(100\,{\rm GeV}/T_{*})^{7/3}&\hbox{helical},\cr 8~{}10^{-14}\,{\rm Gauss}\,(100\,{\rm GeV}/T_{*})^{5/2}&\hbox{non-helical}.\end{cases}$ (2.12)
@@
Fig. 1 shows the primordial magnetic field generated by a supercool phase transition characterized by an efficiency $\epsilon=0.1$, a reheating temperature $T_{*}=\,{\rm TeV}$ and an inverse duration $\beta/H_{*}=30$ corresponding to a characteristic scale $\lambda_{*}H_{*}=(4\pi)^{1/3}/(\beta/H_{*})\approx 0.1$. The dashed curve shows the magnetic field spectrum in the idealised limit where magnetic interactions with the plasma at $a<a_{\rm rec}$ are neglected. Plasma effects are incorporated in the blue and red curves, which correspond to maximal and no helicity, respectively. Since even a small initial helicity tends to grow through inverse cascade, the blue curve likely represents a more realistic outcome. We adopt it for our subsequent plots. Varying the assumed parameters, increasing $\lambda_{*}$ (as possibly favoured by generating Primordial Black Holes as Dark Matter) shifts the spectrum to larger $\lambda_{0}$, while increasing $T_{*}$ reduces both $\lambda_{0}$ and the magnetic field $B_{0}$, preventing accessing the region favoured by blazar observations if $T_{*}\gtrsim 3\,{\rm TeV}$.
@@ 3 Minimal model with two scalar singlets
We are interested in the phase where $\langle s\rangle=w\gg\langle h\rangle=v=246.2$ GeV and $\langle s^{\prime}\rangle=0$. This happens for $R\lesssim 1$ when the critical boundary $\lambda_{HS}=-2\sqrt{\lambda_{H}\lambda_{S}}$ is crossed while $\lambda_{HS}$ is negative. In this phase $s^{\prime}$ is stable. As a result the model contains two DM candidates: the particle $s^{\prime}$ and Primordial Black Holes produced during the phase transition, from regions where bubbles appear later [45].
@@
The colored contours in fig. 2 indicate the peak values of the present-day primordial magnetic field strength, $B_{0}(\lambda_{0})$. Such contours are nearly horizontal because $B_{0}$ is determined as in eq. (2.12) by the reheating temperature, which in the current model is approximated as $T_{*}\approx 0.07m_{s^{\prime}}$. The dot-dashed contours denote the corresponding magnetic correlation length scale $\lambda_{*}H_{*}$. Having assumed a helical magnetic field configuration, sizeable values of both $B_{0}$ and $\lambda_{0}$ — as favored by the blazar anomalies [3, 4, 5] — arise when the reheating temperature $T_{*}$ and the associated new physics scale are not far above the electroweak scale. In contrast, higher values of $T_{*}$ lead to a mostly dark phase transition, suppressing the visible magnetic field signal as discussed in eq. (2.12). Fig. 1 shows an example of the magnetic field spectrum.
The parameter space region potentially favoured by the presence of a sufficiently large primordial magnetic field appears accessible to future experimental tests in multiple ways. First, future dark matter direct detection experiments are expected to reach sensitivities approaching the so-called “neutrino floor”, shown as a dashed curve. Second, the stochastic gravitational wave background produced by the first-order phase transition — computed as in [43] — may be within the reach of future observatories such as LISA and ET. The regions above the dashed curves labeled ‘LISA’ and ‘ET’ correspond to the parameter space where these signals would be detectable with signal-to-noise ratio (SNR) above 10. Finally, gravitational lensing might allow to detect PBH, in case their abundance is not much below the cosmological DM abundance.
@@
Figure 2: The colored contours shows values of the peak value $B_{0}(\lambda_{0})$ of the primordial magnetic field generated by the first order phase transition. The dot-dashed contours show the values of the peak scale $\lambda_{*}H_{*}$, that determines $\lambda_{0}$ as in eq. (2.10). $B_{0}$, computed assuming helical evolution, is intense enough to explain the blazar anomalies below the dashed green curve. The gravitational waves produced by the phase transition are intense enough to be observed at LISA and ET in the regions above the curves denoted with the experiment names. Direct detection experiments exclude the region shaded in red, and their sensitivity will reach the ‘Neutrino Floor’ curve. The magnetic field spectrum of fig. 1 is reproduced at the point denoted with a $\star$.
@@ 4 Model with SU(2)X gauge interaction
Related model of dynamical symmetry contain only two scalars: the Higgs doublet and $S$, such that the scale-invariant scalar potential has the simpler form
@@
In these models the quartics of the scalar $S$ run negative at low energy because it is assumed to be charged under some gauge interaction. We here assume that $S$ is a doublet under an extra SU(2)X with gauge coupling $g_{X}$, such that $\beta_{\lambda_{S}}\simeq{9g_{X}^{4}}/{8}(4\pi)^{2}$. After dynamical symmetry breaking
@@ 5 Model with U(1)B-L gauge interaction
A related models assumes that a complext singlet scalar $S$ with charge $q_{S}$ under gauged U(1)B-L with gauge coupling $g_{B-L}$. Here $B$ is baryon number, $L$ is lepton number, and $B-L$ is their anomaly-free combination. The scalar potential is as in eq. (4.1). The beta function $\beta_{\lambda_{S}}\simeq{6(q_{S}g_{B-L})^{4}}{/(4\pi)^{2}}$ of the quartic $\lambda_{S}$ induces dynamical symmetry breaking at low energy, such that the scalars are expanded as
@@
We again use $m_{s}$ and $M_{Z^{\prime}}$ as the free parameters, where $M_{Z^{\prime}}=q_{S}g_{B-L}w$ is the mass of the $B-L$ vector. In the current model $Z^{\prime}$ decays into quarks and leptons and thereby cannot be stable DM. The gauge coupling, the mixing quartic scalar coupling $\lambda_{HS}$ and the vacuum expectation value $w$ can be written as
@@ 6 Conclusions
We analysed three representative models that realise dynamical symmetry breaking from a classically scale-invariant potential: a minimal model with a second scalar singlet; a model with a hidden SU(2) gauge interaction, a model with a gauged U(1)B-L. These interactions induce radiative corrections that induce dynamical symmetry breaking. For each model, we computed the spectrum of magnetic fields generated during bubble collisions and followed their evolution to the present day, incorporating magnetohydrodynamic effects such as inverse cascades and the role of magnetic helicity. The coherence length of primordial magnetic fields grows in view of the universe expansion and of inverse cascade in the radiation-dominated plasma. This growth depends on the helicity of the initial magnetic field and plasma velocity. As a result, these models can produce magnetic field strengths and coherence lengths consistent with observational hints from blazar-induced $\gamma$-ray spectra, potentially explaining the origin of the inter-galactic magnetic fields. The parameter regions that yield sufficiently intense magnetic fields are also predictive in other ways.
The same first-order phase transitions responsible for magnetogenesis can also generate a stochastic gravitational wave background within reach of upcoming detectors such as LISA and the Einstein Telescope. In addition, these transitions can lead to the formation of primordial black holes (PBHs) in a mass range where they may account for all or part of the dark matter. Mergers of such PBHs produce gravitational wave signals [46]. Some models further include thermal particle dark matter candidates with weak-scale masses and the correct relic abundance from freeze-out, adding another testable signature. In conclusion, these scenarios appear cosmologically testable thanks to their multi-messenger phenomenology.

### v3

@@ 2.1 Magnetic fields at production
where $\Delta V_{H}=m_{h}^{4}/16\lambda_{H}$ with $m_{h}\approx 125\,{\rm GeV}$ and $\lambda_{H}\approx 0.126$ is the energy released by the Higgs part of the potential only, $V_{H}=-m_{h}^{2}h^{2}/4+\lambda_{H}h^{4}/4$; [^1] In theories with an extra singlet scalar $s$, the total energy released can be decomposed as $\Delta V=\Delta V_{s}+\Delta V_{H}$. The energy $\Delta V_{s}$ released from the singlet sector does not source visible magnetic fields. In some models, the singlet is charged under a hidden gauge symmetry, in which case $\Delta V_{s}$ may generate hidden magnetic fields. However, even in the presence of vector kinetic mixing, the contribution of hidden magnetic fields to the visible sector remains negligible [35]. $\epsilon\sim 0.1$ is an efficiency factor for producing magnetic fields [36]; $\kappa$ is given by a sum over physical process leading to magnetic fields. For super-cooled phase transitions with $\alpha\gg 1$, the dominant contribution comes from bubble collisions [^2] Another contribution arises from the detonating expansion of an ultra-relativistic bubble (with velocity $v_{\rm wall}\simeq 1$ for supercooled phase transitions), that transfers to the plasma a fraction $\kappa_{\rm exp}=\frac{\alpha(1-\kappa_{\rm col})^{2}}{0.73+0.083\sqrt{\alpha(1-\kappa_{\rm col})}+\alpha(1-\kappa_{\rm col})}$ (2.2) of the potential energy [37, 18]. The $1-\kappa_{\rm col}$ suppression makes this contribution to $\kappa=\kappa_{\rm col}+\kappa_{\rm exp}$ negligible, at least unless $\alpha$ is huge [18].
@@ 2.2 Magnetic fields today
Instead, during the initial radiation-dominated era, magnetic fields are coupled to the cosmological plasma and undergo inverse cascades. Neglecting the expansion of the universe, the coherence length grows as $\lambda\propto\tau^{p_{\lambda}}$, while the magnetic field strength decreases as $B\propto\tau^{-p_{B}/2}$, where $\tau$ denotes the proper time. Various estimates for the two exponents exist in the literature [41, 42, 39, 43, 44, 40], mainly depending on whether a fraction of the magnetic fields has helical shape, as magnetic helicity $\propto B^{2}\lambda$ is conserved in a plasma with high conductivity: [^4] Magnetic helicity density $\mathbf{A}\cdot\mathbf{B}$ where $\mathbf{B}=\nabla\times\mathbf{A}$ is a measure of the twisting, linking, and knottedness of magnetic field lines. A helical field increases its correlation length as magnetic energy dissipates, driving an inverse cascade of energy from small to large scales.
@@
This means that even a subdominant helical component of the magnetic fields tends to become dominant at late times. Inverse cascades can also arise in the non-helical case, since bubble collisions and parity-violating processes during the phase transition can induce helicity in the plasma velocity field. A more refined approximation may be possible following [44]; however, we do not pursue this direction here due to potential uncertainties related to micro-stability in such estimates. Instead, we rely on the standard approximations. Taking into account the expansion of the universe — where the scale factor evolves as $a\propto\tau$ during the radiation-dominated era — yields
@@
[…]
@@
Fig. 1 shows the primordial magnetic field generated by a supercool phase transition characterized by an efficiency $\epsilon=0.1$, a reheating temperature $T_{*}=\,{\rm TeV}$ and an inverse duration $\beta/H_{*}=30$ corresponding to a characteristic scale $\lambda_{*}H_{*}=(4\pi)^{1/3}/(\beta/H_{*})\approx 0.1$. The dashed curve shows the magnetic field spectrum in the idealised limit where magnetic interactions with the plasma at $a<a_{\rm rec}$ are neglected. Plasma effects are incorporated in the blue and red curves, which correspond to maximal and no helicity, respectively. Since even a small initial helicity tends to grow via inverse cascade, the blue curve likely represents a more realistic outcome. We adopt it for our subsequent plots. Varying the assumed parameters, increasing $\lambda_{*}$ (as possibly favoured by generating Primordial Black Holes as Dark Matter) shifts the spectrum to larger $\lambda_{0}$, while increasing $T_{*}$ reduces both $\lambda_{0}$ and the magnetic field $B_{0}$, preventing accessing the region favoured by blazar observations if $T_{*}\gtrsim 3\,{\rm TeV}$.
@@ 3 Minimal model with two scalar singlets
We are interested in the phase where $\langle s\rangle=w\gg\langle h\rangle=v=246.2$ GeV and $\langle s^{\prime}\rangle=0$. This happens for $R\lesssim 1$ when the critical boundary $\lambda_{HS}=-2\sqrt{\lambda_{H}\lambda_{S}}$ is crossed while $\lambda_{HS}$ is negative. In this phase $s^{\prime}$ is stable. As a result the model contains two DM candidates: the particle $s^{\prime}$ and Primordial Black Holes produced during the phase transition, from regions where bubbles appear later. We employ the estimates from [47], recently questioned in [48, 49].
@@
The colored contours in fig. 2 indicate the peak values of the present-day primordial magnetic field strength, $B_{0}(\lambda_{0})$. Such contours are nearly horizontal because $B_{0}$ is determined as in eq. (2.12) by the reheating temperature, which in the current model is approximated as $T_{*}\approx 0.07m_{s^{\prime}}$. The dot-dashed contours denote the corresponding magnetic correlation length scale $\lambda_{*}H_{*}$. Having assumed a helical magnetic field configuration, size-able values of both $B_{0}$ and $\lambda_{0}$ — as favored by the blazar anomalies [3, 4, 5] — arise when the reheating temperature $T_{*}$ and the associated new physics scale are not far above the electroweak scale. In contrast, higher values of $T_{*}$ lead to a mostly dark phase transition, suppressing the visible magnetic field signal as discussed in eq. (2.12). Fig. 1 shows an example of the magnetic field spectrum.
The parameter space region potentially favoured by the presence of a sufficiently large primordial magnetic field appears accessible to future experimental tests in multiple ways. First, future dark matter direct detection experiments are expected to reach sensitivities approaching the so-called “neutrino floor”, shown as a dashed curve. Second, the stochastic gravitational wave background produced by the first-order phase transition — computed as in [45] — may be within the reach of future observatories such as LISA and ET. The regions above the dashed curves labeled ‘LISA’ and ‘ET’ correspond to the parameter space where these signals would be detectable with signal-to-noise ratio (SNR) above 10. Finally, gravitational lensing might allow detecting PBH, in case their abundance is not much below the cosmological DM abundance.
@@
Figure 2: The colored contours show values of the peak value $B_{0}(\lambda_{0})$ of the primordial magnetic field generated by the first order phase transition. The dot-dashed contours show the values of the peak scale $\lambda_{*}H_{*}$, that determines $\lambda_{0}$ as in eq. (2.10). $B_{0}$, computed assuming helical evolution, is intense enough to explain the blazar anomalies below the dashed green curve. The gravitational waves produced by the phase transition are intense enough to be observed at LISA and ET in the regions above the curves denoted with the experiment names. Direct detection experiments exclude the region shaded in red, and their sensitivity will reach the ‘Neutrino Floor’ curve. The magnetic field spectrum of fig. 1 is reproduced at the point denoted with a $\star$.
@@ 4 Model with SU(2)X gauge interaction
A related model of dynamical symmetry contains two scalars only: the Higgs doublet and $S$, such that the scale-invariant scalar potential has the simpler form
@@
In these models the quartics of the scalar $S$ run negative at low energy because $S$ is assumed to be charged under some gauge interaction. We here assume that $S$ is a doublet under an extra SU(2)X with gauge coupling $g_{X}$, such that $\beta_{\lambda_{S}}\simeq{9g_{X}^{4}}/{8}(4\pi)^{2}$. After dynamical symmetry breaking
@@ 5 Model with U(1)B-L gauge interaction
A related model assumes that a complex singlet scalar $S$ with charge $q_{S}$ under gauged U(1)B-L with gauge coupling $g_{B-L}$. Here $B$ is baryon number, $L$ is lepton number, and BLB-L is their anomaly-free combination. The scalar potential is as in eq. (4.1). The beta function $\beta_{\lambda_{S}}\simeq{6(q_{S}g_{B-L})^{4}}{/(4\pi)^{2}}$ of the quartic $\lambda_{S}$ induces dynamical symmetry breaking at low energy, such that the scalars are expanded as
@@
We again use $m_{s}$ and $M_{Z^{\prime}}$ as the free parameters, where $M_{Z^{\prime}}=q_{S}g_{B-L}w$ is the mass of the BLB-L vector. In the current model $Z^{\prime}$ decays into quarks and leptons and thereby cannot be stable DM. The gauge coupling, the mixing quartic scalar coupling $\lambda_{HS}$ and the vacuum expectation value $w$ can be written as
@@
The scale-invariant BLB-L model has been already studied in [18, 30] at a few benchmark points. Our results agree within a factor of 2 (see e.g. footnote 3) with the benchmark points used there-in, as we assumed the same approximations for the generation of primordial magnetic fields and their cosmological evolution. The agreement holds despite a difference in notations: our efficiency factor $\kappa=1$ corresponds to their $\kappa_{h}=\Delta V_{H}/\Delta V$. Unlike us, these authors also present a second, different, approximation where the full $\Delta V$ (rather than just the smaller jump $\Delta V_{H}$ in the Higgs potential energy) contributes to the magnetic energy density in eq. (2.1). Under this different approximation, that corresponds to $\kappa_{h}=1$ in their notation, the authors of [18, 30] obtain large magnetic fields even if $M_{Z^{\prime}}\gg M_{Z}$. We do not adopt this scenario, as we regard it as unrealistically optimistic, even in the BLB-L model [35].
@@ 6 Conclusions
We analyzed three representative models that realise dynamical symmetry breaking from a classically scale-invariant potential: a minimal model with a second scalar singlet; a model with a hidden SU(2) gauge interaction, a model with a gauged U(1)B-L. These interactions induce radiative corrections that induce dynamical symmetry breaking. For each model, we computed the spectrum of magnetic fields generated during bubble collisions and followed their evolution to the present day, incorporating magnetohydrodynamic effects such as inverse cascades and the role of magnetic helicity. The coherence length of primordial magnetic fields grows in view of the universe expansion and of inverse cascade in the radiation-dominated plasma. This growth depends on the helicity of the initial magnetic field and plasma velocity. As a result, these models can produce magnetic field strengths and coherence lengths consistent with observational hints from blazar-induced $\gamma$-ray spectra, potentially explaining the origin of the inter-galactic magnetic fields. The parameter regions that yield sufficiently intense magnetic fields are also predictive in other ways.
The same first-order phase transitions responsible for magneto-genesis can also generate a stochastic gravitational wave background within reach of upcoming detectors such as LISA and the Einstein Telescope. In addition, these transitions can lead to the formation of primordial black holes (PBHs) in a mass range where they may account for all or part of the dark matter. Mergers of such PBHs produce gravitational wave signals [50]. Some models further include thermal particle dark matter candidates with weak-scale masses and the correct relic abundance from freeze-out, adding another testable signature. In conclusion, these scenarios appear cosmologically testable thanks to their multi-messenger phenomenology.